what is the equation of the line is y=
m=4/5,(6,-8)

Found 2 solutions by algebrahouse.com, mananth:
Answer by algebrahouse.com(1659)

m = 4/5, (6,-8)

Point-slope form is y - y1 = m(x - x1)

y - (-8) = 4/5(x - 6) {substituted into point-slope form}
y + 8 = 4/5(x - 6) {simplified the double negative signs}
y + 8 = 4/5x - 24/5 {used distributive property}
y = 4/5x - 24/5 - 8 {subtracted 8 from both sides}
y = 4/5x - 24/5 - 40/5 {got a common denominator of 5}
y = 4/5x - 64/5 {combined like terms}

For more help from me, visit: www.algebrahouse.com



Answer by mananth(16946)

m= 0.80

Plug value of the slope and point(6,-8) in
Y = m x + b
-8.00 = 4.8 + b
b= -8 - 4.8
b= -12.8
So the equation will be
Y = 0.8x-12.8
y= 4/5 x -64/5
m.ananth@hotmail.ca